OpenStudy (anonymous):

Which value for x makes the sentence true? 9x - 5 = 2x + 16 -2 3 8 5

OpenStudy (anonymous):

\[9x-2x=16+5\]

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

that the clue

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

u wan to know the value of x right

OpenStudy (anonymous):

\[9x-5=2x+16\]